HETAV PATEL

Grade 11,

A ray of light is incident on glass slab r.i- 1.5 thickness t in such manner that the angle of refraction is 60 degree.if the speed of light in vacuum is c then time taken to cross the slab is

A ray of light is incident on glass slab r.i- 1.5 thickness t in such manner that the angle of refraction is 60 degree.if the speed of light in vacuum is c then time taken to cross the slab is

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

If the angle of refraction inside the slab is60^{\circ}, the distance travelled by light inside the slab of thicknesstisd=\dfrac{t}{cos60^{\circ}}=2t
The speed of light in the glass medium=\dfrac{c}{\mu}=\dfrac{2c}{3}
Hence time taken to cross the slab=\dfrac{d}{v}=\dfrac{3t}{c}